OpenStudy (anonymous):

Solve for x in the proportion:

OpenStudy (anonymous):

\[\frac{ 2x }{ 16x + 30 }\] = \[\frac{ -1 }{ x }\]

OpenStudy (johnweldon1993):

What do you get when you do this one?

OpenStudy (anonymous):

2x^2 = -16x - 30 <--- This is all I know to do.

OpenStudy (johnweldon1993):

That is correct.....do me a favor and add 16x and 30 to both sides... You'll then have \[\large 2x^2 + 16x + 30 = 0\] Look familiar?

OpenStudy (anonymous):

So certain problems have to go to = 0? and some don't?

OpenStudy (johnweldon1993):

Well no...but in the form I put it in...it is a quadratic equation....can be solved by factoring...makes it easier than guess and check

OpenStudy (anonymous):

Ahh okay. Let me find the factors of that.

OpenStudy (anonymous):

2(x + 3)(x + 5)

OpenStudy (anonymous):

Then to cancel out to 0, x + 3 would be -3 and x + 5 would be -5? Which is the x's? -3 and -5

OpenStudy (johnweldon1993):

That would be correct! You can even check those answers by plugging then into the original equation

OpenStudy (anonymous):

Alright, thank you. A quick question on a different question: 30x = 18x + 36 you would -18x get 12x = 36 then x = 3 right?

OpenStudy (johnweldon1993):

Correct again!

OpenStudy (anonymous):

Alrighty, thank you sooo much (:

OpenStudy (johnweldon1993):

No problem!